OpenStudy (anonymous):

find all triginometric values of θ with the given conditions tan θ= -1, sin θ < 0 i know its in QIV and that cot θ = -1, but i get stuck there

OpenStudy (anonymous):

I cant think of a way to solve this without having your unit circle memorized. Maybe this might help a little? \[\tan(\theta) = -1 \Longrightarrow \frac {\sin(\theta)}{\cos(\theta)} = -1 \Longrightarrow \sin(\theta) = -\cos(\theta)\] So sin and cos should be the same number, just different signs.

OpenStudy (anonymous):

and your right, its going to be in quadrant IV